1x/2 + 1y/3 = 1/3
1x/6 + 1y/3 = 29

Answer by scott8148(6628)

multiplying 1st eqn by 6 __ 3x+2y=2

multiplying 2nd eqn by 6 __ x+2y=174

subtracting equations __ 2x=-172 __ dividing by 2 __ x=-86

substituting __ (-86)+2y=174 __ adding 86 __ 2y=260 __dividing by 2 __ y=130